OpenStudy (anonymous):

evaluate dy/dt for the function at the point. xy^2=4; dx/dt=-5, x=4, y=1 Answers: 5/8, 8/5, -5/8, -8/5

zepdrix (zepdrix):

So what part are you stuck on? Have you tried taking a derivative yet? Need to apply product rule on the left.

OpenStudy (anonymous):

i've got the derivative of the function, but i don't know what to do to get the dt part.

zepdrix (zepdrix):

We're taking our derivative with respect to t, which means `both` x and y when we take their respective derivatives should give us prime terms, or dx/dt and dy/dt. So taking the derivative gives us: (x)'y^2 + x(y^2)' = (4)' (1)x'y^2 + x(2y)y' = 0 Or if you want it written in Leibniz Notation: (dx/dt)y^2 + 2xy(dy/dt) = 0 Does that look like what you came up with?

OpenStudy (anonymous):

yeah. so now what do I do?

zepdrix (zepdrix):

We have 4 unknowns in that derivative. They provided us with information for 3 of the pieces and we need to solve for the 4th. x = 4 y = 1 dx/dt = -5 Just plug them in and solve for the other one.

OpenStudy (anonymous):

oh, I see! Thank you so much! I've been waiting for an hour, and now I can finally get to bed!

zepdrix (zepdrix):

XD

OpenStudy (anonymous):

I'm serious! I was waiting on one problem for an hour. You saved me another hour or two of waiting. So thank you again.

zepdrix (zepdrix):

np c: go to bed!